Please help me solve this:Write the equation 2x=3y-6 in standard form
Thanks

Answer by Here2Help(19)

Here is what I have:
Initial problem:
2x = 3y + -6
Reorder the terms:
2x = -6 + 3y
Solving
2x = -6 + 3y
Solving for variable 'x'.
Move all terms containing x to the left, all other terms to the right.
Divide each side by '2'.
x = -3 + 1.5y
Simplifying
x = -3 + 1.5y